M. Bot Corporation has 11,000 shares of 8%, $100 par value, cumulative preferred stock outstanding at December 31, 2017. No dividends were declared in 2015 or 2016. If M. Bot wants to pay $410,000 of dividends in 2017, what amount of dividends will common stockholders receive?

Brief Exercise 14-2 M. Bot Corporation has 11,000 shares of 8%, $100 par value, cumulative preferred stock outstanding at December 31, 2017, No dividends were declared in 2015 or 2016. If M. Bot wants to pay $410,000 of dividends in 2017, what amount of dividends w ill common stockholders receive? Amount of dividend LINK TO TEXT
Answer #1

$146,000.

Preffered stock dividend=11,000*100*8%=88,000 per year

Preffered stock dividend payable=88,000*3=$264,000.

amount of dividends will common stockholders receive=410,000-264,000=$146,000.